# Relational Prompt-based Pre-trained Language Models for Social Event Detection

### v1

@@ 5.7. Hyperparameter Study
We explore the impact of five hyperparameters on model performance in the online detection scenario. These hyperparameters include Intra-cluster Loss Weight ($\lambda$), Inter-cluster Loss Weight ($\mu$), Similarity Threshold ($\alpha$), Number of Structured Attention Heads ($o$), and Parameter Update Weight ($\zeta$) during the maintenance phase. To ensure experimental consistency, all parameters except the one under study are kept fixed, with their initial values set to: 0.005, 0.01, 0.9, 2, and 0.4, respectively. Due to computational resource constraints, all hyperparameter experiments are conducted under the $\mathrm{RPLM}_{SED}$ based on RoBERTa-base. The experimental results are shown in Figure 7.
Figure 7. $\mathrm{RPLM}_{SED}$ with different hyperparameters.
@@
Intra-cluster Loss Weight ($\lambda$). When examining the effect of different intra-cluster loss weights on model performance, we select four weights: 0.001, 0.005, 0.01, and 0.015. As shown in Figure 7(a), the model performance remains stable when $\lambda$ is set to 0.001 or 0.005. However, as $\lambda$ increases to 0.01 or higher, a linear decline in model performance is observed. This suggests that an excessively high intra-cluster loss weight causes the model to overly focus on making representations of the same event type more cohesive, neglecting the importance of separating different event types, thereby negatively impacting the final event detection accuracy.
@@
Inter-cluster Loss Weight ($\mu$). We also choose 0.001, 0.005, 0.01, and 0.015 as candidate values for the inter-cluster loss weight. As shown in Figure 7(b), when $\mu$ is set too high (such as 0.015) or too low (such as 0.001), model performance decreases. Especially when $\mu$ is set too low, the model lacks sufficient constraint to push apart the centroids of different clusters, resulting in event clusters not being well separated. Conversely, when $\mu$ is set too high, the model places more emphasis on separating different event clusters, leading to a slight decrease in the cohesiveness of messages within clusters, thus mildly affecting the model’s performance.
@@
Similarity Threshold ($\alpha$). In studying the similarity threshold, we test four thresholds: 0.1, 0.3, 0.6, and 0.9. Figure 7(c) shows that model performance improves with increasing $\alpha$ values. This phenomenon can be attributed to a higher $\alpha$ value implying stricter requirements for message pair similarity. Consequently, the selected candidate representations are more likely to be extracted from message pairs within the same cluster. This leads to more stable and robust aggregated message representations. In addition, the choice of a reasonable similarity threshold is also crucial. Setting $\alpha$ too high may result in message representations that lack comprehensiveness.
@@
Number of Heads in Structured Attention Mechanism ($o$). The experiment examines settings of 1, 2, 3, and 4 attention heads. As illustrated in Figure 7(d), when there is only one attention head, the model is limited in capturing key message features, resulting in limited message representations and consequently affecting model performance. Conversely, too many attention heads may introduce noise, leading to a decline in message representation quality and impacting model performance.
@@
Parameter Update Weight during Model Maintenance ($\zeta$). We select 0.2, 0.4, 0.6, and 0.8 as candidate values for the parameter update weight. As shown in Figure 7(e), the model performs best when $\zeta$ is set to 0.4. When $\zeta$ is set to 0.2, there is a slight decrease in model performance, possibly due to the model retaining too much irrelevant historical information or information unrelated to current events. Conversely, when $\zeta$ values are too high (such as 0.6 and 0.8), the model retains too little historical knowledge, potentially making it difficult for the model to recognize and understand the connection between current and historical events, leading to a performance decline.
Figure 8. Raw Messages of $M_{19}$ within Events2012.
@@ 5.8. Case Study
Firstly, we extract two pairs of messages from $M_{19}$ where FinEvent predicts incorrectly while RPLM predicts accurately. For the ‘case1’ part in Figure 8, $m_{1}$ and $m_{2}$, posted by different users, pertain to positive reactions towards the University of Alabama’s performance in a specific game. The first message provides a broader perspective on the game and mentions a forthcoming challenge, while the second is focused on celebrating a particular scoring moment. Although $m_{1}$ and $m_{2}$ describe the same event, they share no common attributes other than belonging to the same time segment. Unfortunately, FinEvent utilizes time by transforming it into temporal features combined with message content, resulting in no connections between these two messages in the constructed multi-relational message graph. This results in FinEvent’s inability to aggregate and learn effective information about both from the explicit message graph directly, thereby failing to group them correctly. Moreover, in the ‘Case 2’ part from Figure 8, $m_{3}$ articulates an opinion regarding a player, whereas $m_{4}$ delineates a football match. These two messages from different events share the same hashtag ‘#lfc’ and entity ‘Shelvey’, which, in turn, leads to their erroneous classification together by FinEvent. In contrast, $\mathrm{RPLM}_{SED}$ does not rely on the explicit structural relations between messages to learn representations. It transforms the structural relations between messages into multi-relational prompt embeddings to complement message content embeddings, thereby enabling the joint utilization of message content and structural information. During the learning process, by consistently encoding the structural relations and semantic content between messages, it can better learn and understand the complex structural relations between messages and the interaction between structural relations and semantic content. In the absence of common attributes or the presence of misleading information, $\mathrm{RPLM}_{SED}$ can still accurately group them by capturing the semantic and contextual relations between the messages.
@@ 6.1. Social Event Detection
Social event detection represents a pivotal research domain, extensively categorized based on its objectives and application scenarios. In terms of goal-oriented classification, social event detection methodologies are primarily divided into document-pivot (DP) (Cao et al., 2021; Peng et al., 2019; Ren et al., 2022a; Peng et al., 2022; Liu et al., 2020a; Aggarwal and Subbian, 2012; Hu et al., 2017; Zhou and Chen, 2014; Zhang et al., 2007) and feature-pivot (FP) (Fedoryszak et al., 2019; Fung et al., 2005) approaches. DP methods focus on clustering social messages by analyzing their interrelations, emphasizing the grouping of relevant messages for a deeper understanding of the development and evolution of societal events. These methods are particularly crucial in analyzing specific domain events such as disasters, stock trading, major sports events, and political activities, as they aid in organizing and comprehending the historical information of these events. The proposed $\mathrm{RPLM}_{SED}$ method in this paper is an exemplar of this category. In contrast, FP methods cluster messages based on the distribution characteristics of social message elements, such as vocabulary and named entities. They focus on the co-occurrence and distribution patterns of message elements to reveal inter-element correlations. Furthermore, social event detection methodologies can be categorized based on their application scenarios into offline  (Peng et al., 2019) and online methods (Peng et al., 2021; Cao et al., 2021; Ren et al., 2022a; Peng et al., 2022; Liu et al., 2020a; Hu et al., 2017; Fedoryszak et al., 2019; Zhang et al., 2017; Goswami and Kumar, 2016) Offline methods are primarily used for analyzing historical events, whereas online methods continuously monitor social data streams, tracking the real-time development of events. Online methods are crucial for rapid response and real-time event monitoring, especially in situations requiring swift action. Additionally, based on the employed techniques and mechanisms, these methods can be further subdivided into incremental clustering methods (Aggarwal and Subbian, 2012; Hu et al., 2017; Zhang et al., 2007; Ozdikis et al., 2017), community detection methods (Liu et al., 2020a; Fedoryszak et al., 2019; Liu et al., 2018; Yu et al., 2017; Liu et al., 2021, 2020b; Su et al., 2022), and topic modeling methods (Zhou and Chen, 2014; Becker et al., 2011; Cheng et al., 2014; Cordeiro, 2012; Peng et al., 2018; Blei et al., 2003)
@@
Recently, methods based on Graph Neural Networks (GNNs) (Peng et al., 2021; Cao et al., 2021; Cui et al., 2021; Peng et al., 2019, 2022; Ren et al., 2022a, 2021, b, 2021) have demonstrated potential in the domain of social event detection. By modeling social messages as graphs, these approaches effectively integrate messages’ semantic and structural information, overcoming the limitations of insufficient information utilization within traditional methods. For instance, KPGNN (Cao et al., 2021) converts messages with common attributes into homogeneous message graphs, which achieves incremental social event detection by leveraging the inductive learning capabilities of GNNs. FinEvent (Peng et al., 2022) reduces the impact of noisy edges between message nodes by constructing weighted multi-relational graphs and utilizes reinforcement learning to determine optimal relation weights. Additionally, QSGNN (Ren et al., 2022a) employs utilizing the finest known samples and dependable knowledge transfer to extend knowledge from the known to the unknown, markedly diminishing the requirement for manual annotation in model training. However, GNN-based methods learn representations from explicit message graphs. These graphs may contain noisy edges, and messages from the same event without common attributes lack connections. This leads to GNNs being impacted when learning message representations, an effect that is exacerbated in the presence of sparse structural information. Moreover, these methods initialize message embeddings statically before training, preventing the model from dynamically obtaining better message embeddings during the training phases. In recent research, Cao et al. (2024) propose HISEvent, which explores the neighborhood of a message graph by minimizing 1-dimensional Structural Entropy (1D SE) and supplementing edges in the existing message graph between semantically close messages. Subsequently, by minimizing 2-dimensional structural entropy (2D SE) for partitioning the message graph, the approach eliminates the cost of manual annotation, achieving unsupervised social event detection. It is imperative to acknowledge that although HISEvent employs a hierarchical approach to mitigate the computational complexity of conventional vanilla greedy 2D SE minimization algorithm (Li et al., 2016) from $O(N^{3})$ to $O(n^{3})$ (where $N$ and $n$ respectively denote the number of nodes within a message graph and that within a manually selected sub-graph), methods based on GNNs or PLMs exhibit a computational complexity of merely $O(N)$. Given that HISEvent is an unsupervised approach, it has not been compared in this paper. In addition, Our proposed ‘pairwise’ approach significantly differs from that introduced in PPGCN (Peng et al., 2019). In PPGCN, after obtaining representations, the pairwise sampling method selects a positive and a negative sample for each message, addressing the challenge of numerous event types with few messages per event. Unlike PCGNN, our pairwise message modeling strategy is implemented during the data preprocessing stage to preserve structural relations rather than only focusing on inter or intra-event relations.

### v2

@@ 5.8. Hyperparameter Study
We explore the impact of eight hyperparameters on model performance in the online detection scenario. All experiments are conducted on the Events2012 dataset and utilize the HDBSCAN clustering method. These hyperparameters include Intra-cluster Loss Weight ($\lambda$), Inter-cluster Loss Weight ($\mu$), Pairwise Cross-Entropy Loss Weight ($\kappa$), Similarity Threshold ($\alpha$), Number of Structured Attention Heads ($o$), Parameter Update Weight ($\zeta$) during the maintenance phase, Number of message sampled ($y$), and Window Size ($w$). To ensure experimental consistency, all parameters except the one under study are kept fixed, with their initial values set to 0.005, 0.01, 1, 0.9, 2, 0.4, 60, and 3, respectively. Due to computational resource constraints, all hyperparameter experiments are conducted under the $\mathrm{RPLM}_{SED}$ based on RoBERTa-base. The experimental results are shown in Figure LABEL:fig:Hyperparameter.
@@
Intra-cluster Loss Weight ($\lambda$). When examining the effect of different intra-cluster loss weights on model performance, we select four weights: 0.001, 0.005, 0.01, and 0.015. As shown in Figure LABEL:fig:Hyperparameter(a), the model performance remains stable when $\lambda$ is set to 0.001 or 0.005. However, as $\lambda$ increases to 0.01 or higher, a linear decline in model performance is observed. This suggests that an excessively high intra-cluster loss weight causes the model to overly focus on making representations of the same event type more cohesive, neglecting the importance of separating different event types, thereby negatively impacting the final event detection accuracy.
@@
Inter-cluster Loss Weight ($\mu$). We also choose 0.001, 0.005, 0.01, and 0.015 as candidate values for the inter-cluster loss weight. As shown in Figure LABEL:fig:Hyperparameter(b), when $\mu$ is set too high (such as 0.015) or too low (such as 0.001), model performance decreases. Especially when $\mu$ is set too low, the model lacks sufficient constraint to push apart the centroids of different clusters, resulting in event clusters not being well separated. Conversely, when $\mu$ is set too high, the model places more emphasis on separating different event clusters, leading to a slight decrease in the cohesiveness of messages within clusters, thus mildly affecting the model’s performance.
Pairwise Cross-Entropy Loss Weight ($\kappa$). For pairwise cross-entropy loss weight, experiments are conducted on four values: 0.6, 0.8, 1.0, and 1.2. Figure LABEL:fig:Hyperparameter(c) shows a balanced allocation among the weights of pairwise cross-entropy loss, inter-cluster loss, and intra-cluster loss is crucial. The model’s overall performance decreases when the parameter $\kappa$ is set to 0.6 or 0.8. This is attributed to the insufficient constraint imposed by the pairwise cross-entropy loss and a too-small weight, which inversely amplifies the constraints of inter-cluster loss and intra-cluster loss. This may cause similarity scores between messages obtained from the classifier to be inaccurate, thereby affecting the quality of the final message representation. Conversely, an excessively large $\kappa$ value weakens the constraints imposed by both inter-cluster and intra-cluster losses. This, in turn, leads to message representations within the same event not being sufficiently tight and blurring the boundaries between different event clusters, thereby affecting the model’s performance.
@@
Similarity Threshold ($\alpha$). In studying the similarity threshold, we test four thresholds: 0.1, 0.3, 0.6, and 0.9. Figure LABEL:fig:Hyperparameter(d) shows that model performance improves with increasing $\alpha$ values. This phenomenon can be attributed to a higher $\alpha$ value implying stricter requirements for message pair similarity. Consequently, the selected candidate representations are more likely to be extracted from message pairs within the same cluster. This leads to more stable and robust aggregated message representations. In addition, the choice of a reasonable similarity threshold is also crucial. Setting $\alpha$ too high may result in message representations that lack comprehensiveness.
@@
Number of Heads in Structured Attention Mechanism ($o$). The experiment examines settings of 1, 2, 3, and 4 attention heads. As illustrated in Figure LABEL:fig:Hyperparameter(e), when there is only one attention head, the model is limited in capturing key message features, resulting in limited message representations and consequently affecting model performance. Conversely, too many attention heads may introduce noise, leading to a decline in message representation quality and impacting model performance.
@@
Parameter Update Weight during Model Maintenance ($\zeta$). We select 0.2, 0.4, 0.6, and 0.8 as candidate values for the parameter update weight. As shown in Figure LABEL:fig:Hyperparameter(f), the model performs best when $\zeta$ is set to 0.4. When $\zeta$ is set to 0.2, there is a slight decrease in model performance, possibly due to the model retaining too much irrelevant historical information or information unrelated to current events. Conversely, when $\zeta$ values are too high (such as 0.6 and 0.8), the model retains too little historical knowledge, potentially making it difficult for the model to recognize and understand the connection between current and historical events, leading to a performance decline.
Number of message sampled ($y$). We investigate the impact of varying $y$ on model performance, including four different values: 20, 40, 60, and 80. As shown in Figure LABEL:fig:Hyperparameter(g), both excessively large and small $y$ values decrease model performance. For events with a large number of messages, sampling too few may result in insufficiently comprehensive learned message representations. Conversely, social media messages have the long-tail distribution characteristic. When the sampling size is excessively large, the model may overfit the head events with a large number of messages, which neglects the representation learning of messages from long-tail events.
Window Size ($w$). The window size determines the extent of message blocks included during continuous detection and the message block numbers used for model maintenance. We analyze the impact of four different window sizes: 1, 2, 3, and 4. As observed in Figure LABEL:fig:Hyperparameter(h), both excessively small (with w set to 1 or 2) and large window sizes (with w set to 4) result in varying degrees of decline in model performance. Smaller windows are likely to contain more newly emerging or less frequently discussed events. These events’ relevant message counts are relatively less, making the contextual information insufficiently clear. Consequently, during the model maintenance phase, the learned message representations of these events may not be comprehensive. In subsequent window message blocks, when the message numbers describing these events increase, the model struggles to recognize them effectively due to insufficient learning of these events, thereby affecting overall performance. Moreover, the smaller window requires more frequent model maintenance. This may accelerate the forgetting of historical knowledge and lead to the overfitting of current knowledge. Conversely, when the window size is excessively large, some events may have already dissipated in subsequent message blocks. Continuing to fit messages related to these events during model maintenance is unnecessary and may introduce noise, thereby compromising the model’s performance.
@@ 5.9. Case Study
Figure 27. Raw Messages of $M_{19}$ within Events2012.
Firstly, we extract two pairs of messages from $M_{19}$ where FinEvent predicts incorrectly while RPLM predicts accurately. For the ‘case1’ part in Figure 27, $m_{1}$ and $m_{2}$, posted by different users, pertain to positive reactions towards the University of Alabama’s performance in a specific game. The first message is focused on celebrating a particular scoring moment, while the second provides a broader perspective on the game and mentions a forthcoming challenge. Although $m_{1}$ and $m_{2}$ describe the same event, they share no common attributes other than belonging to the same time segment. Unfortunately, FinEvent utilizes time by transforming it into temporal features combined with message content, resulting in no connections between these two messages in the constructed multi-relational message graph. This results in FinEvent’s inability to aggregate and learn effective information about both from the explicit message graph directly, thereby failing to group them correctly. Moreover, in the ‘Case 2’ part from Figure 27, $m_{3}$ articulates an opinion regarding a player, whereas $m_{4}$ delineates a football match. These two messages from different events share the same hashtag ‘#lfc’ and entity ‘Shelvey’, which, in turn, leads to their erroneous classification together by FinEvent. In contrast, $\mathrm{RPLM}_{SED}$ does not rely on the explicit structural relations between messages to learn representations. It transforms the structural relations between messages into multi-relational prompt embeddings to complement message content embeddings, thereby enabling the joint utilization of message content and structural information. During the learning process, by consistently encoding the structural relations and semantic content between messages, it can better learn and understand the complex structural relations between messages and the interaction between structural relations and semantic content. In the absence of common attributes or the presence of misleading information, $\mathrm{RPLM}_{SED}$ can still accurately group them by capturing the semantic and contextual relations between the messages.
@@
5.10. Clustering Algorithm Study
This section investigates the performance of three density-based clustering algorithms, OPTICS (Ankerst et al., 1999), DBSCAN (Ester et al., 1996), and HDBSCAN (Campello et al., 2013), on message representation obtained by $\mathrm{RPLM}_{SED}$. These algorithms are widely used in real-world applications due to their ability to cluster data without requiring a predefined number of event clusters. Specifically, the core of OPTICS is to generate an ordered list of points by analyzing the density and mutual relationships of data points, thus identifying clustering structures at various scales. OPTICS is particularly sensitive to the parameters of Minimum Points and Cluster Reachability. The parameter of Minimum Points determines the minimum number of points in a cluster. Setting this parameter too low may lead to the formation of numerous small clusters, whereas setting it too high may cause small clusters to be merged. Cluster Reachability evaluates the density changes at cluster boundaries to decide whether to split regions into separate clusters. Setting this parameter too low may blur cluster boundaries, causing regions with small density differences to be merged; setting it too high could result in over-segmentation of regions with minor density variations into multiple clusters.
DBSCAN forms clusters by defining core, border, and noise points. It is highly sensitive to the parameters of Epsilon Neighbourhood and Minimum Points. A point is considered a core point if the number of points in its neighborhood is greater than or equal to Minimum Points; otherwise, it is classified as a border or noise point. The algorithm starts from an unvisited core point and identifies all points within its neighborhood to form a new cluster. it recursively applies the same process to the core points within the neighborhood until the cluster can no longer be expanded. When the Epsilon Neighbourhood is set too small, a single cluster may be divided into multiple sub-clusters, instead, set too large, different clusters may be merged. Setting the Minimum Points too low may result in too many core points and ignoring noise points; setting it too high may result in only high-density regions being recognized as clusters, leading to more points being classified as noise.
Compared to DBSCAN, HDBSCAN automatically adjusts the density threshold to recognize clustering structures at varying densities. It constructs a density hierarchy tree to represent data structures at different densities and eventually prunes the tree to obtain the optimal clustering result. The key parameters for HDBSCAN are the Minimum Cluster Size and the Minimum Sample Number. The Minimum Cluster Size specifies the minimum number of points required for a cluster; setting it too low increases the number of noise points, while setting it too high may ignore small clusters. The minimum number of samples determines core points and influences cluster shapes. Typically, the Minimum Cluster Size and the Minimum Sample Number are set to be equal.
Experiments are conducted on the Events2012 dataset in the online detection scenario, in which metrics (NMI, AMI, ARI) for each message block under different algorithms are calculated. The Minimum Points and Cluster Reachability parameters for OPTICS are set to 8 and 0.2, respectively. For DBSCAN, the Epsilon Neighbourhood and Minimum Points are set to 5 and 8, respectively. Additionally, the Minimum Cluster Size for HDBSCAN is set to 8. Figure LABEL:fig:Performance_of_different_clustering_algorithms demonstrates that HDBSCAN is most suitable for the message representations learned by $\mathrm{RPLM}_{SED}$. This is due to HDBSCAN requiring adjustment of only one parameter, the Minimum Cluster Size. In contrast, OPTICS and DBSCAN are more sensitive to multiple parameters. The distribution of data points across different message blocks may vary. Therefore, setting the same Cluster Reachability (OPTICS) or Epsilon Neighborhood (DBSCAN) parameters is not applicable to all message blocks.
@@ 6.1. Social Event Detection
Social event detection represents a pivotal research domain, extensively categorized based on its objectives and application scenarios. In terms of goal-oriented classification, social event detection methods are primarily divided into document-pivot (DP) (Cao et al., 2021; Peng et al., 2019; Ren et al., 2022a; Peng et al., 2022; Liu et al., 2020a; Aggarwal and Subbian, 2012; Hu et al., 2017; Zhou and Chen, 2014; Zhang et al., 2007) and feature-pivot (FP) (Fedoryszak et al., 2019; Fung et al., 2005) approaches. DP methods focus on clustering social messages by analyzing their interrelations, emphasizing the grouping of relevant messages for a deeper understanding of the development and evolution of societal events. These methods are particularly crucial in analyzing specific domain events such as disasters, stock trading, major sports events, and political activities, as they aid in organizing and comprehending the historical information of these events. The proposed $\mathrm{RPLM}_{SED}$ method in this paper is an exemplar of this category. In contrast, FP methods cluster messages based on the distribution characteristics of social message elements, such as vocabulary and named entities. They focus on the co-occurrence and distribution patterns of message elements to reveal inter-element correlations. Furthermore, social event detection methods can be categorized based on their application scenarios into offline  (Peng et al., 2019) and online methods (Peng et al., 2021; Cao et al., 2021; Ren et al., 2022a; Peng et al., 2022; Liu et al., 2020a; Hu et al., 2017; Fedoryszak et al., 2019; Zhang et al., 2017; Goswami and Kumar, 2016). Offline methods are primarily used for analyzing historical events, whereas online methods continuously monitor social data streams, tracking the real-time development of events. Online methods are crucial for rapid response and real-time event monitoring, especially in situations requiring swift action. Additionally, based on the employed techniques and mechanisms, these methods can be further subdivided into incremental clustering methods (Aggarwal and Subbian, 2012; Hu et al., 2017; Zhang et al., 2007; Ozdikis et al., 2017), community detection methods (Liu et al., 2020a; Fedoryszak et al., 2019; Liu et al., 2018; Yu et al., 2017; Liu et al., 2021, 2020b; Su et al., 2022), and topic modeling methods (Zhou and Chen, 2014; Becker et al., 2011; Cheng et al., 2014; Cordeiro, 2012; Peng et al., 2018; Blei et al., 2003)
@@
Recently, methods based on Graph Neural Networks (GNNs) (Peng et al., 2021; Cao et al., 2021; Cui et al., 2021; Peng et al., 2019, 2022; Ren et al., 2022a, 2024, b, 2024; Sun et al., 2023) have demonstrated potential in the domain of social event detection. By modeling social messages as graphs, these approaches effectively integrate messages’ semantic and structural information, overcoming the limitations of insufficient information utilization within traditional methods. For instance, KPGNN (Cao et al., 2021) converts messages with common attributes into homogeneous message graphs, which achieves incremental social event detection by leveraging the inductive learning capabilities of GNNs. FinEvent (Peng et al., 2022) reduces the impact of noisy edges between message nodes by constructing weighted multi-relational graphs and utilizes reinforcement learning to determine optimal relation weights. Additionally, QSGNN (Ren et al., 2022a) employs utilizing the finest known samples and dependable knowledge transfer to extend knowledge from the known to the unknown, markedly diminishing the requirement for manual annotation in model training. However, GNN-based methods learn representations from explicit message graphs. These graphs may contain noisy edges, and messages from the same event without common attributes lack connections. This leads to GNNs being impacted when learning message representations, an effect that is exacerbated in the presence of sparse structural information. Moreover, these methods initialize message embeddings statically before training, preventing the model from dynamically obtaining better message embeddings during the training phases. In recent research, Cao et al. (2024) propose HISEvent, which explores the neighborhood of a message graph by minimizing 1-dimensional Structural Entropy (1D SE) and supplementing edges in the existing message graph between semantically close messages. Subsequently, by minimizing 2-dimensional structural entropy (2D SE) for partitioning the message graph, the approach eliminates the cost of manual annotation, achieving unsupervised social event detection. It is imperative to acknowledge that although HISEvent employs a hierarchical approach to mitigate the computational complexity of conventional vanilla greedy 2D SE minimization algorithm (Li et al., 2016) from $O(N^{3})$ to $O(n^{3})$ (where $N$ and $n$ respectively denote the number of nodes within a message graph and that within a manually selected sub-graph), methods based on GNNs or PLMs exhibit a computational complexity of merely $O(N)$. Given that HISEvent is an unsupervised approach, it has not been compared in this paper. In addition, Our proposed ‘pairwise’ approach significantly differs from that introduced in PPGCN (Peng et al., 2019). In PPGCN, after obtaining representations, the pairwise sampling method selects a positive and a negative sample for each message, addressing the challenge of numerous event types with few messages per event. Unlike PCGNN, our pairwise message modeling strategy is implemented during the data preprocessing stage to preserve structural relations rather than only focusing on inter or intra-event relations.